OpenStudy (anonymous):

Suppose that the functions f and g are defined for all real numbers x as follows f(x)=3x g(x)=4x^2 Write the expressions for (f-g)(x) and (f.g)(x) and evaluate (f+g)(-2)

OpenStudy (primeralph):

@agent0smith Got this?

OpenStudy (agent0smith):

Okay :P f = 3x g= 4x^2 so... f-g = 3x - 4x^2 f.g means f times g, so, f.g = (3x)(4x^2) ...now simplify it f+g = 3x + 4x^2, now find (f+g)(-2) by plugging in x=-2.

OpenStudy (anonymous):

which gives you 10

OpenStudy (agent0smith):

Yep.

OpenStudy (anonymous):

thanks!